OpenStudy (anonymous):

how do i solve log x- log 2 = 2 for x????

OpenStudy (dumbcow):

log x - log 2 = log (x/2)

OpenStudy (theviper):

Tell in brief plz:)

OpenStudy (theviper):

@dumbcow

OpenStudy (anonymous):

it gave me this problem log x - 2= 2 and it says solve for x

OpenStudy (anonymous):

im working with exponential and logarithmic functions

OpenStudy (dumbcow):

i used a log property: log a - log b = log (a/b)

OpenStudy (anonymous):

how do you work it out after u put it in that form

OpenStudy (dumbcow):

then you have log (x/2) = 2 convert to exponential form x/2 = 10^2 x/2 = 100 x = 200

OpenStudy (anonymous):

correct thank you

OpenStudy (theviper):

k!k! I understood Thanx:)

OpenStudy (anonymous):

do u guys know how to work this other one out 2 log x - 3 log (y+1) i have to put it in single logarithm